OpenStudy (anonymous):

Id greatly appreciate it if you could answer this question! hydrated nickel (II) chloride is a green crystalline compund. when heated strongly, the compound is dehydrated. if 0.23 g of NiCl2. xH2O gives 0.128 g of NiCl2 on heating what is the value of x?

OpenStudy (jfraser):

in a chemical formula, the numbers refer to the relative number of molecules, not their masses, so you have to find the number of molecules (or moles) of water that are lost by heating, and the number of molecules (or moles) of nickel II chloride left behind when the heating is done.\[0.23g nickel hydrate - 0.128g nickel salt = 0.102g water "lost"\]so\[0.102g H_2O * \frac{1mol H_2O}{18g H_2O} = 0.0057mol H_2O "lost"\] and \[0.128g NiCl_2 * \frac{1mol NiCl_2}{129.8g NiCl_2} = 0.00099 mol NiCl_2\]Now these really small moles can't go in a formula, but the RATIO of these number can. Divide each one by the smaller of the 2, and you're guaranteed to turn 1 of them into a small whole number.\[\frac{0.0057mol H_2O}{0.00099mol NiCl_2} = \frac{5.75mol H_2O}{1mol NiCl_2} \approx 6\] the value of x is approximately 6.

OpenStudy (anonymous):

thank you so much!